# version 1.0

### v6

@@ 4.3 Controllability of Linear Dynamical Systems
Recall that a system is controllable if and only if in the absence of noise and for every target state there is a sequence of controls that drives the system state to that target. Linear dynamical systems exhibit a particularly elegant characterization of controllability. First, we need to define an object of central importance for linear dynamical systems, the Kalman matrix.
@@
Consider the evolution of the noiseless linear dynamical system starting from the initial state $\mathbf{x}_{0}$. The state at time $T$ is given by the expression
@@
By the fundamental theory of linear algebra, this equation has a solution if and only if ${K}_{T}({A,B})$ has full rank. That is, for every initial state $\mathbf{x}_{0}$ and final state $\mathbf{x}_{T}$, there exists a set of control vectors $\{\mathbf{u}_{t}\}$ that satisfy the equation if and only if ${K}_{T}({A,B})$ has full rank. The rank of the Kalman matrix is upper bounded by the number of rows of ${K}_{T}$, which is $d_{x}$. Thus, the system is controllable if and only if the rank of ${K}_{T}$ is $d_{x}$ for all $T\geq d_{x}$.
Denote by $r_{t}=\mathop{\mbox{\rm rank}}({K}_{t})$ the rank of the Kalman matrix after $t$ iterations. Since the rank can never decrease with $t$, it remains to show that the system is controllable if and only if $r_{t}=r_{d_{x}}$ for all $t\geq d_{x}$.
We claim that once $r_{t}=r_{t+1}$, the rank never increases again by increasing $t$. Assuming $r_{t}=r_{t+1}$, we have
$\mathop{\mbox{\rm span}}[{B,AB,\ldots,A^{t}B}]=\mathop{\mbox{\rm span}}[{B,AB,\ldots,A^{t-1}B}].$
Since linear transformations retain containment, we can multiply by A{A} and get
$\mathop{\mbox{\rm span}}[{AB,\ldots,A^{t+1}B}]\subseteq\mathop{\mbox{\rm span}}[{AB,\ldots,A^{t}B}].$
Adding B{B} to both matrices does not affect the subset relationship.
$\mathop{\mbox{\rm span}}[{B,AB,\ldots,A^{t+1}B}]\subseteq\mathop{\mbox{\rm span}}[{B,AB,\ldots,A^{t}B}]$
Therefore, $r_{t+2}\leq r_{t+1}$. Since the rank cannot decrease through the addition of columns, we have $r_{t+2}=r_{t+1}$. Since the rank can grow at most $d_{x}$ times, equation $r_{d_{x}}=r_{T}$ holds for all $T>d_{x}$ if and only if the system is controllable, and this concludes the theorem. ∎
@@ 4.4 Quantitative Definitions
In the rest of this text our goal is to give efficient algorithms with finite time performance bounds for online control. These finite-time bounds naturally depend on how difficult it is for a given system to stabilize. Therefore, we require a quantitative definition for stabilizability, one that holds for both time-invariant and time-varying linear dynamical systems.
@@ 4.4.1 Stabilizability of Time-Invariant Systems
Consider a linear time-invariant system given by the transformations (A,B)({A,B}). We have already remarked that this system is called stabilizable, according to the classical definition, if and only if there exists $K$ such that $\rho({A+BK})<1$ holds.
@@
Consider a linear controller K{K} for the LTI system (A,B)({A,B}). If there exist $\kappa>0$ and $\delta\in(0,1]$ such that the matrix A+BK{A+BK} admits the following inequality for all $t$, then $K$ is a $\gamma$-stabilizing controller for $(A,B)$.
$\|(A+BK)^{t}\|\leq\kappa(1-\delta)^{t},\qquad\frac{\kappa}{\delta}\leq\gamma$
@@
Let $\tilde{K}=A+BK$}. By unrolling the dynamics equation we have that
$\mathbf{x}_{t+1}$ $={A}\mathbf{x}_{t}+{B}\mathbf{u}_{t}+\mathbf{w}_{t}=\tilde{K}\mathbf{x}_{t}+\mathbf{w}_{t},$
$=\sum_{i=0}^{t-1}\tilde{{K}}^{i}\mathbf{w}_{t-i}+\tilde{{K}}^{t+1}\mathbf{x}_{0}.$
Thus, as long as $\|\mathbf{w}_{t}\|,\|\mathbf{x}_{0}\|\leq 1$, we have
$\|\mathbf{x}_{t+1}\|$ $\leq\sum_{i=0}^{t}\|\tilde{K}^{i}\|\|\mathbf{w}_{t-i}\|+\|\tilde{K}^{t+1}\|\|\mathbf{x}_{0}\|,$
$\leq\kappa\sum_{i=0}^{\infty}(1-\delta)^{i}\leq\frac{\kappa}{\delta}\leq\gamma.$
@@
4.4.2 Stabilizability of Time-Varying Systems
The $\gamma$-stabilizability condition, as opposed to classical stabilizability, is particularly important to generalize to time-varying systems, since the spectral radius of the product is not bounded by the product of spectral radii, see exercise 4.6 at the end of this chapter.
We now define a sufficient condition for $\gamma$-stabilizability. Consider a time-varying linear dynamical system given by the transformations $\{{A}_{t},{B}_{t}\}$.
Lemma 4.9.
Consider a sequence of linear controllers ${K}_{1:T}$ for the time-varying linear dynamical system $\{{A}_{1:T},{B}_{1:T}\}$. If the matrices $\tilde{K}_{t}=A_{t}+B_{t}K_{t}$} satisfy the following inequality, then this linear controller sequence is $\gamma$-stabilizing.
$\forall t_{1},t_{2}\ .\ \left\|\prod_{t=t_{1}}^{t_{2}}\tilde{K}_{t}\right\|\leq\kappa(1-\delta)^{t_{1}-t_{2}},\ \frac{\kappa}{\delta}\leq\gamma.$
Proof.
Notice that under the condition of the lemma, by unrolling the dynamics equation we have that
$\|\mathbf{x}_{t+1}\|$ $=\|{A}_{t}\mathbf{x}_{t}+{B}_{t}\mathbf{u}_{t}+\mathbf{w}_{t}\|=\left\|\tilde{{K}}_{t}\mathbf{x}_{t}+\mathbf{w}_{t}\right\|,$
$=\left\|\sum_{i=0}^{t-1}(\prod_{j=1}^{i}\tilde{{K}}_{t-j})\mathbf{w}_{t-i}+(\prod_{j=1}^{t}\tilde{{K}}_{t-j})\mathbf{x}_{0}\right\|,$
$\leq\kappa\sum_{i=0}^{\infty}e^{-\delta i}\leq\frac{\kappa}{\delta}\leq\gamma.$
@@ 4.4.3 Controllability of Time-Invariant Systems
Clearly, such a system is nominally controllable, as is indeed any one-dimensional system with nonzero $B$. However, reaching a state $x=1$ from an initial $x_{0}=0$ in a constant number of time steps requires a large control input $u_{t}$ of magnitude at least $1/\varepsilon$. Thus, such a system may be said to be barely controllable. This difficulty of controllability is captured in the definition below, which requires not only that $K(A,B)$ is a full row-rank, but also that such a statement is robust to small changes in $A,B$.
@@
$\mathop{\mbox{\rm rank}}{K}_{k}({A},{B})=d_{x},\text{ and }\|({K}_{k}(A,B){K}_{k}(A,B)^{\top})^{-1}\|\leq\kappa.$
In the above definition, in addition to the full row-rank constraint, the spectral norm of the inverse of the Gram matrix of $K(A,B)$ is bounded by $\kappa$. For example, in the simple one-dimensional system we presented above, with $k=1$, the matrix $K(A,B)$ is $\varepsilon$, and thus $\kappa=\frac{1}{\varepsilon^{2}}$, which indeed captures the complexity of controlling the system.
@@ 4.5 Bibliographic Remarks
The idea of reducing a nonlinear dynamical system to a time varying linear dynamical system had been independently discovered in disparate fields. In the context of control, using the linearization to characterize local stability properties of non-linear systems was pioneered by Lyapunov [Lya92], and is called the Lyapunov direct method; see, e.g. [SL91].
The asymptotic notions of stability, stabilizability and controllability discussed in this chapter constitute a core part of theory of linear control, and are found in most texts on the subject (e.g. [Ste94, Hes18]).
The efficient verification of stabilizability via a semidefinite program and the recovery of a stable controller form the solution of the latter appears in [Ahm16]. On the other hand, it is known that even imposition of norm constraints on a plausibly stable controller makes this problem computationally hard [BT97, BT00, BT99]. Somewhat distinct but related problems of stabilizability with even a 1-dimensional linear constraint are open to the date.
The notion of $\gamma$-stabilizability has a precursor in the notion of strong stability introduced in [CHK+18]. An extension of this strong stability concept to time-varying linear dynamics was established in [CKM19]. The definition presented in this chapter has weaker requirements than those in the listed articles, allowing us to work with and guarantee regret later with respect to a larger class of comparator policies.
@@ 4.6 Exercises
4.2. Prove that $\rho(A)<1$ if and only if $\lim_{t\mapsto\infty}A^{t}=0$. Furthermore, if $\rho(A)>1$, then for any norm, $\lim_{t\mapsto\infty}\|A^{t}\|=\infty$. Finally, prove that $\rho(A)=\lim_{t\to\infty}\|A^{t}\|^{1/t}$.
@@
$\rho(\prod_{i=1}^{n}A_{i})\nleq\prod_{i=1}^{n}\rho(A_{i}).$
4.4. Prove that a stochastic linear dynamical system with Gaussian perturbations
@@
Hint: One way to do this exercise is to explicitly calculate the stationary distribution of an LDS and then to verify reversibility as stated in its definition. Exercise 3.6 provides an alternative without the need to derive the stationary distribution.
@@
A. For a time-invariant linear system $(A,B)$ equipped with a linear policy $K$, prove that BIBO stability indeed implies $\rho(A+BK)<1$. In fact, comment on how large $\rho(A+BK)$ can be.
B. For a time-invariant BIBO-stable linear system, prove that there exist $\kappa>0$ and $\delta\in(0,1]$ such that, for all $t$, $\|(A+BK)^{t}\|\leq\kappa(1-\delta)^{t}$. In particular, this proves that the conditions listed in Lemma 4.8 are necessary, in addition to being sufficient.
C. Construct a time-varying linear system that is BIBO stable but does not have the exponential decay properties required in Lemma 4.9.
4.6. Prove that the bound of one on the operator norm of the matrix $A$ is a sufficient but not necessary condition for stabilizability.
@@
In the chapters thus far, we have studied more basic notions than control, namely solution concepts of dynamical systems such as stabilizability and controllability. Even for these more basic problems, we have seen that they are computationally intractable in general.
This motivated us to look into simpler dynamics, namely linear dynamical systems, for which we can at least answer basic questions about stabilizability and controllability. In this chapter, we consider how we can control linear dynamical systems optimally.
The answer is positive in a strong sense. An archetypal problem in optimal control theory is the control of linear dynamical systems with quadratic cost functions. This setting is called the Linear Quadratic Regulator which we define and study in this chapter. The resulting efficient and practical algorithm is one of the most widely used subroutines in control.
@@ 5.1 The Linear Quadratic Regulator
The linear quadratic regulator (LQR) problem is that of optimal control of a linear dynamical system with known and fixed quadratic costs. For simplicity, we mostly consider the time-invariant version, where the linear dynamical system is fixed, and the costs are also fixed quadratic, as given formally below:
@@
The (time-invariant) Linear Quadratic Regulator (LQR) input is a linear dynamical system that evolves according to
$\mathbf{x}_{t+1}={A}\mathbf{x}_{t}+{B}\mathbf{u}_{t}+\mathbf{w}_{t},$
where $\mathbf{w}_{t}\sim\mathcal{N}(0,\sigma^{2})$ is i.i.d. Gaussian, and convex quadratic cost functions are specified by matrices $Q$ and $R$.
The goal is to generate a sequence of controls $\mathbf{u}_{t}$ that minimizes the expected cost, taken over the randomness of the perturbations, given by quadratic loss functions,
$\min_{\mathbf{u}(\mathbf{x})}\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{x}_{t}^{\top}Q\mathbf{x}_{t}+\mathbf{u}_{t}^{\top}\mathbf{R}\mathbf{u}_{t}\right].$
@@
    $\mathbf{w}_{t}\in\mathbb{R}^{d_{w}}$ is the perturbation sequence drawn i.i.d. from normal distributions, or any other zero-mean distribution as long as the draws are independent of states or controls.
@@
    $A\in{\mathbb{R}}^{d_{x}\times d_{x}}$, $B\in{\mathbb{R}}^{d_{x}\times d_{u}}$ are the system matrices.
@@
    $Q\in{\mathbb{R}}^{d_{x}\times d_{x}}$, ${R}\in{\mathbb{R}}^{d_{u}\times d_{u}}$ are positive definite matrices. Thus, the cost functions are convex.
@@
The LQR is a special case of a Markov decision process which we have described in Chapter 3, since the transitions are stochastic. However, we cannot naively apply value iteration, policy iteration, or linear programming since there are infinitely many states and actions.

### v8

@@ 4.3 Controllability of Linear Dynamical Systems
Recall that a system is controllable if and only if, in the absence of noise and for every target state, there is a sequence of controls that drives the system state to that target. Linear dynamical systems exhibit a particularly elegant characterization of controllability. First, we define an object of central importance for linear dynamical systems: the Kalman controllability matrix.
@@
Consider the evolution of the noiseless linear dynamical system starting from the initial state $\mathbf{x}_{0}$. The state at time $T$ is given by
@@
Therefore, for every initial state $\mathbf{x}_{0}$ and target state $\mathbf{x}_{T}$, there exists a control sequence $\mathbf{u}_{0},\ldots,\mathbf{u}_{T-1}$ that satisfies this equation if and only if the column span of ${K}_{T}(A,B)$ is all of ${\mathbb{R}}^{d_{x}}$, i.e., if and only if ${K}_{T}(A,B)$ has rank $d_{x}$.
Denote by $r_{t}=\mathop{\mbox{\rm rank}}({K}_{t})$ the rank of the Kalman matrix after $t$ steps. Since adding columns cannot decrease rank, the sequence $(r_{t})_{t\geq 1}$ is nondecreasing. We claim that once $r_{t}=r_{t+1}$, the rank can never increase again.
Indeed, if $r_{t}=r_{t+1}$, then
$A^{t}B\in\mathop{\mbox{\rm span}}[B,AB,\ldots,A^{t-1}B].$
Multiplying by $A$ gives
$A^{t+1}B\in\mathop{\mbox{\rm span}}[AB,A^{2}B,\ldots,A^{t}B]\subseteq\mathop{\mbox{\rm span}}[B,AB,\ldots,A^{t}B].$
Hence adding the next block $A^{t+1}B$ does not increase the rank, so $r_{t+2}=r_{t+1}$. By induction, once the rank stops increasing, it never increases again.
Since the rank can increase at most $d_{x}$ times, it follows that $r_{t}=r_{d_{x}}$ for all $t\geq d_{x}$. Therefore, the system is controllable if and only if
$\mathop{\mbox{\rm rank}}K_{d_{x}}(A,B)=d_{x}.$
@@ 4.4 Quantitative Definitions
In the rest of this text our goal is to give efficient algorithms with finite-time performance bounds for online control. These bounds naturally depend on how difficult it is to stabilize and control a given system. Since the main focus of the sequel is on linear time-invariant systems, we now introduce quantitative notions of stabilizability and controllability for that setting.
@@ 4.4.1 Stabilizability of Time-Invariant Systems
Consider a linear time-invariant system given by the pair (A,B)({A},{B}). We have already remarked that this system is classically stabilizable if and only if there exists a matrix $K$ such that $\rho(A+BK)<1$.
@@
Consider a linear controller K{K} for the LTI system (A,B)({A},{B}). If there exist $\kappa>0$ and $\delta\in(0,1]$ such that
$\|(A+BK)^{t}\|\leq\kappa(1-\delta)^{t}\qquad\text{for all }t\geq 0,$
and moreover $\kappa/\delta\leq\gamma$, then $K$ is a $\gamma$-stabilizing controller for $(A,B)$.
@@
Let $\tilde{K}=A+BK$}. Under the linear controller $u_{t}=Kx_{t}$, the dynamics are
$\mathbf{x}_{t+1}=\tilde{K}\mathbf{x}_{t}+\mathbf{w}_{t}.$
Unrolling the recursion gives
$\mathbf{x}_{t+1}=\tilde{K}^{t+1}\mathbf{x}_{0}+\sum_{s=0}^{t}\tilde{K}^{t-s}\mathbf{w}_{s}.$
Define an augmented disturbance sequence by $v_{-1}=\mathbf{x}_{0}$ and $v_{s}=\mathbf{w}_{s}$ for $s\geq 0$. Then
$\mathbf{x}_{t+1}=\sum_{s=-1}^{t}\tilde{K}^{\,t-s}v_{s}.$
Therefore, whenever $\|\mathbf{x}_{0}\|\leq 1$ and $\|\mathbf{w}_{s}\|\leq 1$ for all $s$,
$\|\mathbf{x}_{t+1}\|$ $\leq\sum_{s=-1}^{t}\left\|\tilde{K}^{\,t-s}\right\|\|v_{s}\|$
$\leq\kappa\sum_{i=0}^{t+1}(1-\delta)^{i}$
$\leq\frac{\kappa}{\delta}\leq\gamma.$
Thus $K$ is $\gamma$-stabilizing. ∎
@@ 4.4.2 Controllability of Time-Invariant Systems
Clearly, such a system is nominally controllable, as is indeed any one-dimensional system with nonzero $B$. However, reaching a state $x=1$ from an initial state $x_{0}=0$ in a constant number of time steps requires a large control input $u_{t}$ of magnitude at least $1/\varepsilon$. Thus, such a system may be said to be barely controllable. This difficulty is captured by the definition below, which requires not only that $K_{k}(A,B)$ has full row rank, but also that the corresponding Gram matrix be well conditioned.
@@
$\mathop{\mbox{\rm rank}}{K}_{k}({A},{B})=d_{x},\qquad\left\|({K}_{k}(A,B){K}_{k}(A,B)^{\top})^{-1}\right\|\leq\kappa.$
In the above definition, in addition to the full-row-rank constraint, the spectral norm of the inverse of the Gram matrix of $K_{k}(A,B)$ is bounded by $\kappa$. For example, in the simple one-dimensional system above with $k=1$, the matrix $K_{1}(A,B)$ is just $\varepsilon$, and thus $\kappa=1/\varepsilon^{2}$, which indeed captures the complexity of controlling the system.
@@ 4.5 Bibliographic Remarks
The idea of reducing a nonlinear dynamical system to a time-varying linear dynamical system has appeared independently in several fields. In the context of control, using linearization to characterize local stability properties of nonlinear systems was pioneered by Lyapunov [LYA92], and is called the Lyapunov direct method; see, e.g., [SL91].
The asymptotic notions of stability, stabilizability, and controllability discussed in this chapter constitute a core part of linear control theory, and are found in most texts on the subject (e.g. [STE94, HES18]).
The efficient verification of stabilizability via a semidefinite program, together with the recovery of a stabilizing controller from a solution to that program, appears in [AHM16]. On the other hand, it is known that even imposing norm constraints on a plausibly stable controller makes this problem computationally hard [BT97, BT00, BT99]. Somewhat distinct but related problems of stabilizability with even a 1-dimensional linear constraint remain open to this day.
The notion of $\gamma$-stabilizability has a precursor in the notion of strong stability introduced in [CHK+18]. An extension of this strong stability concept to time-varying linear dynamics was established in [CKM19]. The definition presented in this chapter has weaker requirements than those in the cited articles, allowing us to work with and guarantee regret later with respect to a larger class of comparator policies.
@@ 4.6 Exercises
4.2. Prove that
$\rho(A)<1\quad\text{if and only if}\quad\lim_{t\to\infty}A^{t}=0.$
Furthermore, if $\rho(A)>1$, then for any norm,
$\lim_{t\to\infty}\|A^{t}\|=\infty.$
Finally, prove that
$\rho(A)=\lim_{t\to\infty}\|A^{t}\|^{1/t}.$
@@ 4.2. Prove that
$\rho\!\left(\prod_{i=1}^{n}A_{i}\right)\nleq\prod_{i=1}^{n}\rho(A_{i}).$
4.4. Assume $\rho(A)<1$. Prove that the stochastic linear dynamical system with Gaussian perturbations
@@
Hint: One way to do this exercise is to explicitly calculate the stationary distribution of the chain and then verify detailed balance directly.
@@
A. For a time-invariant linear system $(A,B)$ equipped with a linear policy $K$, prove that BIBO stability implies $\rho(A+BK)<1$. In fact, comment on how large $\rho(A+BK)$ can be.
B. For a time-invariant BIBO-stable linear system, prove that there exist $\kappa>0$ and $\delta\in(0,1]$ such that, for all $t$,
$\|(A+BK)^{t}\|\leq\kappa(1-\delta)^{t}.$
In particular, this proves that the conditions listed in Lemma 4.6 are necessary, in addition to being sufficient.
4.6. Prove that the strict condition $\|A\|<1$ on the operator norm of $A$ is sufficient, but not necessary, for stabilizability.
@@
In the chapters thus far, we have studied more basic notions than optimal control, namely structural properties of dynamical systems such as stabilizability and controllability. Even for these more basic problems, we have seen that they are computationally intractable in general.
This motivated us to look into simpler dynamics, namely linear dynamical systems, for which we can at least answer basic questions about stabilizability and controllability. In this chapter, we consider how to control linear dynamical systems optimally.
The answer is positive in a strong sense. An archetypal problem in optimal control theory is the control of linear dynamical systems with quadratic cost functions. This setting is called the Linear Quadratic Regulator, which we define and study in this chapter. The resulting efficient and practical algorithm is one of the most widely used subroutines in control.
@@ 5.1 The Linear Quadratic Regulator
The linear quadratic regulator (LQR) problem is that of optimal control of a linear dynamical system with known and fixed quadratic costs. For simplicity, we mostly consider the time-invariant version, where the dynamics and the costs do not vary with time.
@@
The finite-horizon (time-invariant) Linear Quadratic Regulator (LQR) problem consists of a linear dynamical system evolving according to
$\mathbf{x}_{t+1}=A\mathbf{x}_{t}+B\mathbf{u}_{t}+\mathbf{w}_{t},\qquad t=1,\dots,T,$
where the perturbations $\mathbf{w}_{t}\in{\mathbb{R}}^{d_{x}}$ are i.i.d. with
$\mathbb{E}[\mathbf{w}_{t}]=0,\qquad\mathbb{E}[\mathbf{w}_{t}\mathbf{w}_{t}^{\top}]=\Sigma.$
Given positive definite matrices $Q\in{\mathbb{R}}^{d_{x}\times d_{x}}$ and $R\in{\mathbb{R}}^{d_{u}\times d_{u}}$, the goal is to choose policies $\pi_{1},\dots,\pi_{T}$, where $\mathbf{u}_{t}=\pi_{t}(\mathbf{x}_{t})$, that minimize the expected quadratic cost
$\min_{\pi_{1},\dots,\pi_{T}}\mathbb{E}\!\left[\sum_{t=1}^{T}\mathbf{x}_{t}^{\top}Q\mathbf{x}_{t}+\mathbf{u}_{t}^{\top}R\mathbf{u}_{t}\right].$
@@
    $\mathbf{w}_{t}\in{\mathbb{R}}^{d_{x}}$ is the perturbation sequence, assumed i.i.d. with zero mean and covariance $\Sigma$. A Gaussian disturbance is a common special case, but Gaussianity is not essential for the derivation below.
@@
    $A\in{\mathbb{R}}^{d_{x}\times d_{x}}$ and $B\in{\mathbb{R}}^{d_{x}\times d_{u}}$ are the system matrices.
@@
    $Q\in{\mathbb{R}}^{d_{x}\times d_{x}}$ and $R\in{\mathbb{R}}^{d_{u}\times d_{u}}$ are positive definite matrices. Thus, the stage costs are strictly convex in the control.
@@
The LQR is a special case of a Markov decision process, which we described in Chapter 3, since the transitions are stochastic. However, we cannot naively apply value iteration, policy iteration, or linear programming since there are infinitely many states and actions.